# NormAd: A Benchmark for Measuring the Cultural Adaptability of Large Language Models

### v5

@@ 3.2 Story Filtration
After filtration, we have 2633 stories across covering all 75 countries and labels. Detailed statistics across each cultural bin from the Inglehart-Welzel cultural map are provided in Table 2 in Appendix A.1. We include individual prompts used for the various checks in Appendix A.2.
3.3 Human Story Validation
To validate these stories after filtration, we sample a set of 480 stories equally distributed across all 3 ground-truth labels (Yes, No, Neutral), with each story accompanied by a Rule-of-Thumb contextualization. For human validation, we deliberately exclude Value and Country contexts to focus solely on validating the stories under the comprehensive Rule-of-Thumb. Further, its unrealistic for human annotators to adequately validate stories under the Value and Country settings without access to external knowledge. They may also encounter potential cultural biases stemming from a limited understanding of social norms across different cultures. We task two graduate student volunteers for this story validation. We find that our two annotators have a very high agreement (Cohen’s $\kappa=0.95$). Additionally, the majority vote of the two annotators for the RoT setup (with ties broken arbitrarily) shows a very high Cohen’s $\kappa$ agreement of $0.934$ with our gold labels. We compute the RoT accuracy using majority voting and report a 95.6% overall accuracy with ground truth labels, which serves as a measure of human performance. Label-wise accuracies are 96% for ‘Yes’, 92% for ‘No’, and 98% for ‘Neutral’.
Next, to validate of the generated Rule-of-Thumb and Value, we conduct two more human validation studies for the entailment Checks 2 and 3 from §3.2. We perform this over a subset of 144 examples, and report very high agreements between humans and GPT-4 ($\kappa=1,0.86$ respectively).
Having achieved high agreement rates between human and GPT-4 evaluators, we conclude that no further modifications on NormAd are required. Our benchmark, NormAd, is generated by heavily relying on rigorously validated cultural knowledge from Cultural Atlas as context and utilizes GPT-4 as a medium to weave culturally diverse norms into stories, ensuring that GPT-4’s inherent biases or knowledge gaps do not infiltrate our dataset. The high agreement rates from our human validation experiments affirm this.
@@ 3.3 Human Story Validation
(c) Value
(d) Country
Figure 3: Comparison of accuracies across LLaMa-1-SFT (7b, 13b, 30b), LLaMa-2 (7b, 13b, 70b), OLMo7b (SFT/Chat), GPT-3.5-turbo, GPT-4, and Mistral over the all three contexts. Models perform significantly worse in Country and Value contexts compared to the RoT context. Baseline performance (no context) is reported in Figure 11 in Appendix B
@@ 4 Experimental Setup
Models
We utilize NormAd to assess the cultural adaptability of current models, spanning open-source and closed-source LLMs. The models evaluated encompass a wide scope, differing in the number of parameters and finetuning objectives.
Setup and Metrics
@@
5 Results
We evaluate several models on NormAd and analyze across different dimensions.
5.1 RQ1: How well do models perform across different levels of cultural contexts?
@@
We report the accuracies for Country and Value contexts, as shown in Figure 3, suggesting that LLMs struggle under these settings. Top models, like GPT-3.5-turbo, GPT-4footnote 5 and Mistral-7b-Instruct, manage only around 60% accuracy for Value and 51-55% for Country. Ideally, LLMs should demonstrate better adaptability, especially to Value, which relates to abstract, higher-level principles of a norm. Incorporating both Country and Value into the context (Cval; Figure 3(b)) results in slight performance improvements. Mistral-7b-Instruct and Llama-2-70b-chat models fare better than the standalone Country context. The GPT-based models, Mistral-7b-Instruct and the smaller variants of Llama-2 and OLMo also show marginal gains over mere Value contexts.
@@
Evaluating the social acceptability of a story under a specific RoT, which contains all the necessary information to navigate the specific situation, the QA essentially simplifies into a task of textual similarity or entailment. Through our story validation (3.3), we see that humans achieve a high accuracy of 95.6% on this task. However, models seem to underperform, potentially due to a lack of adaptability to textual similarity under cultural and social nuances. The best performing models are GPT-4555 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance. with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance. These findings highlight the lack of cultural adaptability in LLMs, and thereby necessitate further work in this direction.
@@
We observe in Figure 3 that model performance improves with increasing number of parameters (though not linearly), as demonstrated by Llama-2-chat (7b, 13b, 70b) and Llama-1 (supervised finetuned SFT for 7b, 13b, 30b) with regards to RoT and Value conditioning. However, it is a little unclear why the largest models (Llama-2-70b-chat and Llama-1-30b) underperform significantly under Country.
5.2 RQ2: How well do models perform across the Inglehart-Welzel (IW) cultural map?
We have mapped 75 countries from our dataset into 8 clusters based on the Inglehart-Welzel cultural map. The Value conditioned results, illustrated in Figure 5, show that top-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 5 dramatically vary in performance across different cultural zones. For instance, they perform better with narratives from “English Speaking” countries like the USA, but lesser so with “African-Islamic” countries such as Saudi Arabia. Conversely, poorer-performing models, like Llama-2-7b and Llama-1-30b-SFT, have consistently low performance across all cultural zones. Our hypothesis is that as model size increases or training regime improves, models may be better at identifying and exploiting western cultural cues, resulting in an increased skewed performance distribution across cultural zones of the world. We see similar trends across RoT and Country (see Appendix B.2). This ‘western-centric’ bias is commensurate with model performance over other datasets Johnson et al. (2022); Naous et al. (2023) and has been shown across different LM architectures Palta and Rudinger (2023) and even across modalities Ventura et al. (2023).
@@ 5.2 RQ2: How well do models perform across the Inglehart-Welzel (IW) cultural map?
Figure 4: Effect of preference alignment over the accuracies of LLaMa-1 models, against the Value context. KTO improves performance significantly for 30b parameter models, with lesser improvement for 7b models.
@@
Recent training regimes involving Reinforcement Learning from Human Feedback (RLHF) claim to enable LLMs, trained on a general corpus of text data, to align to complex human values (Ziegler et al., 2019; Stiennon et al., 2020; Glaese et al., 2022; Bai et al., 2022; Ouyang et al., 2022). We investigate the effect of different optimization methods used for RLHF on the cultural adaptability of LLMs. We evaluate PPO (Offline) (Schulman et al., 2017), DPO (Rafailov et al., 2024) and KTO (Ethayarajh et al., 2024) on supervised finetuned (SFT) Llama 1 models  [^6] Archangel suite from ContextualAI. We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger Llama-1-30b model. Figure 4 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on Value. We see similar trends with respect to RoT and Country as well (see App. B.1).
Figure 5: Model-wise accuracies across the African-Islamic and English Speaking cultural zones under Value. Top-performing models show a notable performance skew, performing better on stories from English-speaking countries.
5.3 RQ3: What is the performance across subcategories of Etiquette?
@@
We analyse model performance across the 4 subcategories: ‘Eating’, ‘Gifting’, ‘Visiting’, and ‘Basic Etiquette’. We find that models consistently underperform on the ‘Gifting’ subcategory, even when contextualized with the Rule-of-Thumb. In contrast, ‘Eating’, which models find easiest under RoT, along with ‘Visiting’ and ‘Basic Etiquette’, show improved performance with RoT conditioning. After a qualitative analysis, we see that stories under ‘Gifting’ involve highly detailed norms, which include the presentation, number, and color of gifts. Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and the significance playing a significant role in societal norms (Hanna and Srivastava, 2015). This highlights the considerable adaptability required from LLMs in handling such complex social norms. Table 4 in Appendix B.4 presents some failure cases of GPT-3.5-turbo in the ‘Gifting’ and ‘Eating’ subcategories.
5.4 RQ3: How well do models do across social acceptability in story (Yes/No/Neutral)?
Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 5) across ground-truth labels. Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e. violations) and “neutral” (i.e irrelevant).
@@ 5.4 RQ3: How well do models do across social acceptability in story (Yes/No/Neutral)?
We analyze how the social acceptability of stories affects model performance. Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 5). Models generally perform better on stories that conform to social norms (labeled ‘Yes’). For stories that violate social norms (labeled ‘No’), we observe a performance improvement with increasing levels of context. This trend suggests that the inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
@@
Interestingly, most models show a degradation in performance when neither social adherence nor violation is present in the story (labeled ‘Neutral’), even under Rule-of-Thumb contexts (accuracy = $0.42$). This suggests model overconfidence compared to humans, who achieve 98% accuracy for the neutral label (§3.3). The varied performance across social acceptabilities highlights the need to address LLMs’ agreement or sycophancy biases to improve cultural adaptability as also shown in Sun et al. (2024); Perez et al. (2022).
6 Discussion
On ‘culture’ and NormAd
Defining ‘culture’ is challenging due to its multifaceted nature. Adilazuarda et al. (2024) acknowledge this and offer a taxonomy categorizing related work by culture-proxies, linguistic interactions, and measurement strategies. Using their taxonomy, our work can be categorized as a black-boxed LM-evaluation, using a semantic proxy of culture by testing models’ reasoning ability over social norms and etiquette (§3.1.1), with analyses on demographically informed axes (§5.2). We examine the objectives and values of norms and how LLMs reason with them through our Value setup. However, our main distinction is on the task itself, by measuring cultural reasoning by varying the contextualizations provided to the model.
On the adaptability of LLMs
There is a growing consensus among researchers that LLMs need better contextualization capabilities (Richardson, 2021; Fortuna et al., 2022; Yerukola et al., 2023). We show that current LLMs struggle with adhering to and adapting to the diverse and evolving cultural nuances across the world. While LLMs should encode a better foundational understanding of cultures worldwide, they should also be equipped with specific mechanisms to handle the social norms provided in context. We propose prioritizing adaptability at inference time rather than attempting to encode all cultural knowledge into the models’ parameters. Further, we should strive to collect more holistic preference datasets that incorporate demographic or cultural identity to help imbue cultural adaptability into LLMs during finetuning.
@@ 7 Conclusion
In this work, we curate NormAd, a dataset of 2.6k stories involving everyday social etiquette situations spanning across 75 countries. Each story instantiates a cultural norm of a country, distilled into varying levels of cultural context: a fine-grained RoT, an abstracted Value and Country name. Additionally, the stories in NormAd are equally distributed across four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gifting’, and are labeled based on adherence to social norms (‘Yes’, ‘No’, ‘Neutral’). We find that models struggle across the different levels of contexts, with Value and Country posing the most challenges. While larger models and the KTO optimization method improve performance, this leads to increased performance skew across cultural zones, with English-speaking countries performing the best. Models particularly struggle with stories in the ‘Gifting’ subcategory which involves adhering to presentation, number and color of gifts. Further, they also exhibit inherent sycophancy biases, performing significantly better on stories conforming to social norms than those that involve a violation or a neutral scenario. These findings underscore the need for better contextualization capabilities, and highlight the need for more nuanced cultural adaptability in LLMs.
Limitations
Our research examines the cultural reasoning capabilities exhibited by large language models. However, it is imperative to acknowledge certain limitations within our study regarding cultural adaptability, which may warrant further exploration:
Large scale manual evaluation: Our dataset is grounded in an extensively verified source, i.e. the cultural atlas. Additionally, we conduct extensive human evaluations across all contexts to determine dataset validity. However, in an ideal scenario, a human evaluation study over NormAd would additionally necessitate a large-scale evaluation with annotators over diverse backgrounds and cultures, demanding an unbiased human effort. Such a large scale effort involving annotators over all 75 countries would prove challenging owing to the significant resources constraints and effort required.
Fine-Grained Cultural Norms: While our curated test bed encompasses narratives from 75 countries, it is essential to note the existence of more than 75 countries globally, each characterized by unique and finely-grained cultural norms. Additionally, there can be multiple cultural variations within a single country. We opted to scope our study based on the Cultural Atlas as a verified source of information, recognizing that certain nuances may not be fully captured within this framework.

### v6

@@ 4.2 Automatic Filtration
After filtration, we have 2633 stories across covering all 75 countries and 3 labels. Detailed statistics across each cultural bin from the Inglehart-Welzel cultural map are provided in Table 2 in Appendix A.5.
4.3 Human Validation
Validation of gpt-4 Filtration
To validate the filtration proxy of gpt-4 in §4.2, we randomly sampled a subset of 144 data points across 8 Inglehart-Welzel clusters, 4 subcategories, and 3 labels countries (1-2 per label). Two graduate students (Indian demographic) manually verified the quality and validity of the generated RoT and Value. We observed a very high agreement between the human evaluations and gpt-4 for both checks, with Cohen’s $\kappa=1.0,0.86$ respectively.
Dataset Validation
We additionally conduct human validation using Amazon Mechanical Turk (MTurk). For cost reasons, we randomly sample 300 data points stratified across 75 countries, 4 subcategories, and 3 labels (1 data point per label). We qualify annotators from USA, Mexico and India. Each data point is validated by 3 workers. For each data point, we ask workers to perform five subtasks:
  1. 1.
    RoT $\xleftrightarrow{}$ Cultural Atlas Verify that the RoT is derived from the provided country-specific social norms (from Cultural Atlas).
  2. 2.
    Value $\xleftrightarrow{}$ RoT Confirm that the Value is a relevant high-level abstraction of the RoT.
  3. 3.
    Value $\xleftrightarrow{}$ Cultural Atlas Ensure that the Value is relevant to the provided country-specific social norms.
  4. 4.
    Situation $\xleftrightarrow{}$ RoT Verify that the situation is relevant to and revolves around the RoT.
  5. 5.
    Label $\xleftrightarrow{}$ Situation + RoT Finally, given the situation description and the RoT, check if the gold label (Yes/No/Neutral) is correct.
Annotators endorsed our checks’ validity at 84.2% on average, and their interrater agreement yielded a Fleiss fixed-marginal multirater $\kappa=0.56$ and pairwise agreement (PPA) $=0.73$. These results indicate that the annotators overwhelmingly rated our situations, corresponding RoT s, Values, labels, and their relationships as valid, confirming the validity of NormAd. We report per-question scores and payments in Appendix B
@@ 4.3 Human Validation
(c) Country
Figure 3: Comparison of accuracies across LLaMa-1-SFT (7b, 13b, 30b), LLaMa-2 (7b, 13b, 70b), OLMo7b (SFT/Chat), GPT-3.5-turbo, GPT-4, and Mistral over the all three contexts. Models perform significantly worse in Country and Country +Value contexts compared to the RoT context. Baseline performance (no context) is reported in Appendix C and D.
4.4 Verification of Human Performance
We ask humans to determine the most appropriate label for a situation, mimicking the model evaluation setup (unlike §4.3 which involves verifying the gold label). We consider two setups:
Situational Description + RoT
For this setup, we sample 480 data points, stratified across 4 subcategories, 3 labels, and 8 Inglehart-Welzel cultural bins, ensuring at least three data points in each group. We employ 2 graduate students (Indian demographic) for this. We find a very high agreement between the annotators, with Cohen’s $\kappa=0.95$. We compute the RoT accuracy through majority voting (breaking ties arbitrarily), reporting an overall accuracy of 95.6% against the ground truth labels. The label-wise accuracies are 96% for ‘Yes’, 92% for ‘No’, and 98% for ‘Neutral’. This showcases that humans have a strong ability to judge the acceptability of situations when provided with fine-grained RoT contexts.
Situational Description + Value + Country
We conduct a small-scale human study considering 3 countries: India, China, South Korea. For each country, we sample 12 data points across 4 subcategories, and select 1 data point per label. We employ 3 graduate students from each country. Averaging across 3 countries, we achieve a Krippendorff’s $\alpha=0.45$ and Fleiss’s $\kappa=0.63$. We compute Value + Country accuracy through majority voting, reporting an overall accuracy of 91.6% against the ground truth labels. The label-wise accuracies are 90% for ‘Yes’, 86.7% for ‘No’, and 91.6% for ‘Neutral’. This highlights that humans from the relevant culture show strong performance at determining the acceptability of situations when conditioned on abstract Value and Country contexts. See App. B.3 for country-wise splits.
@@ 5 Experimental Setup
5.1 Models
We utilize NormAd-Eti to assess the cultural adaptability of current models, spanning open-source and closed-source LLMs. The models evaluated encompass a wide scope, differing in the number of parameters and finetuning objectives.
5.2 Setup and Metrics
@@
6 LLM Culture Adaptability Results
We evaluate several models on NormAd-Eti and analyze across different dimensions.
6.1 How well do models perform across different levels of cultural contexts?
@@
LLMs show clear limitations when handling Country and Value +Country contexts, with the best performing models GPT-3.5-turbo, GPT-4footnote 6, and Mistral-7b-Instruct achieving only 59-63% accuracy for Value +Country and 51-56% for Country (see Figure 3). In contrast, our human study across three countries (§4.4) demonstrates that humans can perform very well in these settings, achieving a high accuracy of 90%. The wide performance gap highlights the pressing need for LLMs to better adapt to Country and Value contexts, given that real-world scenarios might often lack specificity wrt cultural cues.
@@
Evaluating the social acceptability under RoT is straightforward since it contains all the necessary information to navigate the specific situation. The QA task essentially reduces to a contextual textual similarity or entailment problem. Our human study (§4.4) demonstrates that humans perform exceptionally well on this task, achieving high 95.6% accuracy. However, models under perform, likely due to a lack of adaptability to textual similarity under cultural and social nuances. The best performing models are GPT-4666 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance. with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance. These findings highlight the gap in contextualization capabilities of LLMs, especially with respect to cultural contexts.
@@
We observe in Figure 3 that model performance improves with increasing number of parameters (though not linearly), as demonstrated by Llama-2-chat (7b, 13b, 70b) and Llama-1 (supervised finetuned SFT for 7b, 13b, 30b) with regards to RoT context. The largest models (Llama-2-70b-chat and Llama-1-30b) likely underperform with the Country context, possibly due to insufficient context for eliciting appropriate cultural responses Mukherjee et al. (2024).
6.2 How well do models perform across the Inglehart-Welzel (IW) cultural map?
We mapped 75 countries into 8 clusters based on the Inglehart-Welzel cultural map. The Country + Value conditioned results, illustrated in Figure 4, show that best-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 6 vary in performance across different cultural zones. For instance, they perform better with situations from “English Speaking” countries (e.g., USA) than from “African-Islamic” countries (e.g., Saudi Arabia). In contrast, poorer-performing models, like Llama-2-13b and Llama-1-30b-SFT, under perform consistently across all zones. We hypothesize that larger model sizes and improved training regimes lead to better exploitation of Western cultural cues, causing skewed performance across zones. We see similar trends across Country and RoT (see Appendix D.2). This ‘western-centric’ bias is consistent with model performance on other datasets Johnson et al. (2022); Naous et al. (2023) and observed across various LM architectures Palta and Rudinger (2023) and modalities Ventura et al. (2023).
Figure 4: Model-wise accuracies under Country + Value show a notable performance skew, with top models performing better in social situations from English-speaking countries than in African-Islamic cultural zones.
@@ 6.2 How well do models perform across the Inglehart-Welzel (IW) cultural map?
Recent training regimes involving Reinforcement Learning from Human Feedback (RLHF) claim to enable LLMs, trained on a general text data, to align with complex human values (Ziegler et al., 2019; Stiennon et al., 2020; Glaese et al., 2022; Bai et al., 2022; Ouyang et al., 2022). We investigate the impact of different optimization methods – PPO (Offline) (Schulman et al., 2017), DPO (Rafailov et al., 2024) and KTO (Ethayarajh et al., 2024) – on cultural adaptability of LLMs, specifically focusing on supervised finetuned (SFT) Llama 1 models  [^7] Archangel suite from ContextualAI.
Figure 5: Effect of preference alignment over the accuracies of LLaMa-1 models, against the RoT context. KTO improves performance significantly for 30b parameter models, with lesser improvement for 7b models.
We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger 30b model. Figure 5 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on RoT. We see similar trends for Country and Value + Country as well (see App. D.1).
6.3 What is the performance across subcategories of NormAd-Eti?
@@
We analyse model performance across the 4 subcategories: ‘Eating’, ‘Gifting’, ‘Visiting’, and ‘Basic Etiquette’. Models consistently underperform in ‘Gifting’, even with RoT conditioning, while they excel in ‘Eating’ and show improved results in ‘Visiting’ and ‘Basic Etiquette’. Our qualitative analysis reveals that ‘Gifting’ involves highly detailed norms regarding the presentation, number, and color of gifts. Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and meaning playing a significant role in societal norms (Hanna and Srivastava, 2015). This highlights the considerable adaptability required from LLMs in handling such complex social norms. Table 6 in App. D.4 presents some failure cases of GPT-3.5-turbo.
6.4 How well do models do across social acceptabilities (Yes/No/Neutral)?
@@
We analyze how the social acceptability labels of situations affects model performance. Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6, Llama-1-30b-SFT-KTO). Models generally perform better on situations that conform to social norms (labeled ‘Yes’), and improve on norm-violating situations (labeled ‘No’) with increasing levels of context, indicating that inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
@@
Interestingly, most models show performance degradation when neither social adherence nor violation occurs in social situations (labeled ‘Neutral’), achieving only 42% accuracy even under RoT. This indicates a potential overconfidence in the models, as humans achieve 98% accuracy for neutral labels (§4.4). The varied performance across social acceptabilities highlights the need to address LLMs’ agreement or sycophancy biases to improve cultural adaptability as also shown in Sun et al. (2024); Perez et al. (2022).
Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6) across ground-truth labels. Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e. violations) and “neutral” (i.e irrelevant).
@@ 7 Conclusion
In this work, we introduce a novel hierarchical evaluation framework, NormAd, to assess the contextual adaptability of LLMs, a departure from most prior work which only probes cultural knowledge. Instantiating this framework, we constructed NormAd-Eti, a dataset of 2.6k social etiquette related situations spanning across 75 countries, evaluated for varying degrees of cultural norm specificity: specific Country names, abstract high-level Values with Country names, and fine-grained RoT. Further, NormAd-Eti involves four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gifting’, with three labels of adherence to social norms (‘Yes’, ‘No’, ‘Neutral’).
@@
We find that models struggle across all levels of contexts, particularly with Country +Value and Country setups, lagging significantly behind human performance. While larger models and KTO optimization perform better, we see an increased performance skew across cultural zones, with English-speaking countries performing the best. Models face significant challenges in the ‘Gifting’ subcategory, which involves adhering to presentation, number, and color of gifts. Further, they also exhibit inherent sycophancy biases, performing significantly better on situations conforming to social norms. These findings underscore the need for better contextualization, and more nuanced cultural adaptability in LLMs.
8 Limitations
Our research examines LLM’s ability to adapt to cultural nuances through a test bed of social situations. However, certain limitations present in our dataset and evaluation framework may warrant further research, such as:
Existence of Multiple Cultural Proxies:
Adilazuarda et al. (2024) acknowledges the difficulty of defining what ‘culture’ means in the context of language models, and categorize prior work according to cultural proxies, linguistic interactions and measurement strategies. NormAd can be thus categorized as a black-box LM evaluation employing a semantic proxy to test models’ reasoning abilities concerning etiquette-related social norms (§4.1), with analyses on demographically informed axes (§6.2). While our approach provides valuable insights on LLM’s cultural competencies, we believe that a more comprehensive picture of LLMs’ cultural landscape can be painted through other demographically and lingusitcally informed proxies and frameworks, and , by broadening the scope of evaluation Bhatt and Diaz (2024).
Multilingualism and Linguistic Variations: